# Restore Text First, Enhance Image Later: Two-Stage Scene Text Image Super-Resolution with Glyph Structure Guidance

### v1

@@ Appendix A Appendix
A.1 LLM Acknowledgments
@@
A.2 Dataset Collection
We use VIVO X200 Ultra to collect images for 4 separate focal lengths (14 mm, 35 mm, 84 mm, and 200 mm). We first use PP-OCRV5 Cui et al. (2025) for the rough annotation, then we manually filter the images and annotations. During the filtering and annotation, each image undergoes the following rules:
@@ A.2 Dataset Collection
A.3 More Implementation Details
@@
Dataset Settings. For training, we combine a synthetic dataset with real paired datasets (Real-CE (Ma et al., 2023) and UltraZoom-ST). Our synthetic dataset builds upon LSDIR (Li et al., 2023b), containing 27,000 triplets $(x_{H},x_{L},x_{m})$. We render text on the GT of LSDIR and the corresponding text mask using the LBTS (Tang et al., 2023), then apply Real-ESRGAN degradation Wang et al. (2021) to generate LR images. Since there are misaligned image pairs in Real-CE (Hu et al., 2025; Zhang et al., 2024a) and text lines not annotated, we filter out misaligned image pairs and reannotate the images manually. In the end, we obtain 337 training pairs and 188 testing pairs from the Real-CE dataset. Images under 13mm and 52mm focal lengths are considered as LR $x_{L}$ and GT $x_{H}$, respectively. Following Hu et al. (2025), we use SAM-TS (Ye et al., 2024) to obtain the text mask $x_{m}$ from $x_{H}$. UltraZoom-ST triplets $(x_{H},x_{L},x_{m})$ are processed identically. For stage 1 training, we use cropped text regions from the dataset, while for stage 2, we use full images. Following Tuo et al. (2024), we use PP-OCRv3 Li et al. (2022) to extract strings from the text region of predicted SR for evaluation of OCR-A.
@@
Training Details. For stage 1, we build our model based on the IDM baseline of DiffTSR (Zhang et al., 2024a). We set $\lambda_{td}$, $\lambda_{Seg}$, $\lambda_{Focal}$, and $\lambda_{Dice}$ as $1$, $0.1$, 2020, and $1$ respectively. In stage 2, our model is built upon Stable Diffusion 3.5 medium (Esser et al., 2024) and follows a similar tile-based inference strategy to TADiSR (Hu et al., 2025) and SUPIR (Yu et al., 2024). The timestep $t$ is set to $150$. $\lambda_{l2}$, $\lambda_{LPIPS}$, and $\lambda_{edge}$ are set to $1$, 55, 100100 when using the synthetic dataset for pretraining. Then, $\lambda_{edge}$ is set to 0 and the remaining coefficients are kept unchanged when training on real paired datasets. We train both stages of the model using the AdamW (Loshchilov & Hutter, 2017) optimizer and set the learning rate to $5\times 10^{-5}$ and $5\times 10^{-6}$ for stages 1 and 2 separately. All experiments are conducted on NVIDIA H20 GPUs. For stage 1, we train the model with synthetic and real datasets for 8 epochs, and then use synthetic data to finetune the model for 2 epochs. For stage 2, we first pretrain the model on the synthetic dataset for 50 epochs. Then we train on the real paired datasets for 50 epochs.
A.4 Additional results
We provide detailed quantitative results of UltraZoom-ST in Table 6. Additional qualitative results are provided in Fig. 6.
A.5 Static and Examples of UltraZoom-ST
@@
In Table 7, we provide detailed statistics on the composition of the UltraZoom-ST dataset. Additionally, in Fig. 7, we present some example images from the dataset.
A.6 Efficiency Analysis
@@
As shown in Table 8, stage 1 is a standard diffusion process that takes multiple steps in inference, the efficiency of our model may be suboptimal compared to one-step methods. However, our method achieves state-of-the-art performance in OCR-A, which cannot be easily obtained by extending inference time.
A.7 Discussions and Limitations
Our method leverages an OCR model to localize and interpret text in low-resolution images. In cases of severe degradation, however, OCR fails to recognize the text, which leads to failures in our method as well. Human-provided annotations could potentially mitigate this limitation.
A further challenge arises in real-world deployment. Because our text restoration model is diffusion-based, it requires multiple inference steps, which hinders real-time performance. Although our model reconstructs text structures more plausibly than prior approaches, it does not yet satisfy the speed demands of practical applications. Strategies such as model distillation and quantization offer promising directions to address this bottleneck.
This work is presented as an academic study of scene text image super-resolution in real-world conditions. Still, given its relevance to many applications, it carries both potential benefits (e.g., enhancing image quality in consumer devices) and risks (e.g., exposing private information from photos on social media). On balance, the societal benefits far outweigh the risks, particularly since complementary techniques such as inpainting and watermarking can be used to safeguard sensitive information.
@@ A.7 Discussions and Limitations
Figure 7: Detailed examples of UltraZoom-ST.

### v2

@@ 6 Conclusion
Table 10: Ablation on the effect of stochastic OCR outputs on the performance of our method. The performance gains as the randomness of the OCR output drops. Even under 100% random OCR output, our method still outperforms TADiSR, proving low reliance on OCR.
[…]
TADiSR 24.68 0.795 0.362 0.227 52.13 35.5%
TIGER (Ours) performance under different randomness of OCR output.
100% Random Text 25.60 0.833 0.188 0.150 28.92 40.3%
50% Random Text 25.60 0.833 0.189 0.151 28.92 40.6%
20% Random Text 25.63 0.834 0.185 0.150 29.3 43.7%
0% Random Text (Ours) 25.63 0.834 0.185 0.150 28.82 44.6%
[…]
85mm ($\times 2.35$) Real-ESRGAN 25.84 0.854 0.139 0.130 28.25 23.59 0.843 0.170 0.168 60.8%
HAT 27.33 0.873 0.128 0.133 27.21 24.76 0.870 0.139 0.146 62.0%
MACRONet 22.90 0.804 0.234 0.173 40.19 17.92 0.693 0.515 0.360 54.2%
SeeSR 25.02 0.826 0.167 0.159 32.31 21.62 0.796 0.245 0.207 28.2%
SupIR 25.58 0.823 0.202 0.160 29.33 21.91 0.808 0.239 0.214 33.3%
DiffTSR 23.48 0.805 0.230 0.157 35.77 17.92 0.688 0.392 0.285 48.2%
DiffBIR 25.25 0.793 0.175 0.152 24.57 22.26 0.803 0.218 0.190 36.8%
OSEDiff 26.38 0.852 0.153 0.137 23.92 22.84 0.833 0.200 0.204 41.0%
DreamClear 25.31 0.811 0.169 0.151 24.19 21.44 0.790 0.253 0.221 26.8%
TSD-SR 23.99 0.800 0.161 0.165 30.40 20.62 0.786 0.232 0.218 37.8%
DiT4SR 24.71 0.810 0.173 0.143 29.15 21.25 0.800 0.221 0.193 48.2%
TADiSR 26.31 0.858 0.177 0.149 30.47 23.55 0.856 0.175 0.178 55.7%
Ours 27.00 0.871 0.123 0.120 22.34 24.31 0.860 0.135 0.142 63.2%
35mm ($\times 5.71$) Real-ESRGAN 23.70 0.785 0.218 0.184 42.62 20.86 0.782 0.272 0.235 34.1%
HAT 25.05 0.819 0.207 0.182 41.32 21.98 0.816 0.250 0.241 34.6%
MACRONet 22.28 0.774 0.284 0.199 49.14 17.43 0.694 0.523 0.361 31.5%
SeeSR 23.76 0.795 0.198 0.175 37.92 20.55 0.778 0.271 0.222 25.2%
SupIR 23.49 0.746 0.302 0.204 41.80 20.00 0.757 0.335 0.282 24.0%
DiffTSR 22.57 0.773 0.281 0.184 45.78 16.99 0.671 0.443 0.312 31.2%
DiffBIR 23.76 0.731 0.245 0.196 35.99 20.82 0.770 0.275 0.234 28.6%
OSEDiff 25.35 0.827 0.192 0.166 29.06 21.62 0.809 0.263 0.263 29.9%
DreamClear 24.20 0.778 0.209 0.177 32.08 20.66 0.779 0.283 0.251 23.9%
TSD-SR 22.93 0.757 0.200 0.199 37.61 19.64 0.760 0.282 0.246 27.3%
DiT4SR 23.45 0.774 0.204 0.158 31.73 20.12 0.779 0.250 0.208 26.5%
TADiSR 24.68 0.795 0.362 0.227 52.13 21.57 0.798 0.366 0.347 35.5%
Ours 25.63 0.834 0.185 0.150 28.82 22.14 0.823 0.204 0.194 44.6%
14mm ($\times 14.29$) Real-ESRGAN 22.09 0.718 0.423 0.284 89.22 18.87 0.721 0.503 0.427 11.5%
HAT 22.66 0.738 0.451 0.299 89.58 19.27 0.738 0.533 0.484 12.2%
MACRONet 21.00 0.716 0.422 0.254 79.67 17.01 0.680 0.553 0.382 10.1%
SeeSR 21.79 0.731 0.310 0.228 64.31 18.87 0.717 0.396 0.308 13.2%
SupIR 21.38 0.681 0.447 0.268 70.09 17.99 0.690 0.487 0.383 11.0%
DiffTSR 20.88 0.713 0.412 0.233 74.07 15.87 0.630 0.538 0.348 12.0%
DiffBIR 21.61 0.630 0.392 0.255 69.64 18.61 0.685 0.438 0.328 11.3%
OSEDiff 23.11 0.768 0.274 0.214 49.76 19.44 0.740 0.385 0.334 12.7%
DreamClear 22.45 0.719 0.364 0.260 62.95 19.14 0.725 0.440 0.377 11.7%
TSD-SR 21.14 0.705 0.273 0.225 57.07 17.67 0.686 0.393 0.293 11.9%
DiT4SR 20.88 0.707 0.282 0.181 47.45 17.36 0.687 0.367 0.253 11.7%
TADiSR 22.43 0.721 0.555 0.320 98.44 19.20 0.730 0.585 0.518 14.4%
Ours 23.41 0.774 0.301 0.209 54.74 19.73 0.748 0.374 0.322 16.0%
Total RealEsrGAN 23.99 0.790 0.248 0.194 30.60 21.25 0.786 0.302 0.266 37.1%
HAT 25.17 0.815 0.249 0.198 30.12 22.18 0.813 0.291 0.276 37.9%
MARCONet 22.13 0.768 0.306 0.205 34.28 17.48 0.690 0.529 0.366 33.4%
SeeSR 23.64 0.788 0.219 0.184 26.73 20.45 0.767 0.297 0.241 22.8%
SupIR 23.62 0.754 0.308 0.207 27.91 20.10 0.756 0.344 0.287 23.6%
DiffTSR 22.41 0.767 0.300 0.189 31.25 17.00 0.665 0.452 0.313 31.7%
DiffBIR 23.67 0.724 0.262 0.197 23.10 20.70 0.757 0.302 0.245 26.6%
OSEDiff 25.07 0.819 0.201 0.169 20.53 21.43 0.798 0.276 0.263 28.9%
DreamClear 24.10 0.773 0.238 0.191 21.75 20.50 0.768 0.317 0.276 21.4%
TSD-SR 22.79 0.757 0.207 0.194 24.08 19.42 0.748 0.296 0.249 26.6%
DiT4SR 23.16 0.767 0.215 0.159 20.58 19.73 0.760 0.273 0.216 23.7%
TADiSR 24.61 0.796 0.203 0.160 36.61 21.59 0.799 0.360 0.336 36.6%
Ours 25.48 0.830 0.196 0.156 20.01 22.22 0.814 0.228 0.212 43.0%
7 Reproducibility Statement
To ensure reproducibility, we have made the following efforts:
  1. 1.
    We will release our code and dataset.
  2. 2.
    We provide implementation details in Sec. 5.1 and Sec. 11, including the training process and selection of hyper-parameters.
  3. 3.
    We provide details on evaluation metrics and dataset preparation in Sec. 4 and Sec. 10, and the code and data will be made available along with it.
8 Ethics Statement
This work focuses on improving scene text image super-resolution to support beneficial applications such as enhancing accessibility, document restoration, and navigation assistance. However, we acknowledge potential risks, including misuse in privacy-sensitive contexts (e.g., recovering text from personal or social media images) or unintended deployment in surveillance. To mitigate such risks, users are encouraged to combine our methods with privacy-preserving techniques such as watermarking or selective inpainting. Our UZ-ST dataset was collected from public, non-sensitive scenes under ethical guidelines, without including personally identifiable or private information. We believe the societal benefits of improved text image restoration outweigh potential risks, provided that the technology is applied responsibly.
9 LLM Acknowledgments
@@
10 Dataset Collection
We use VIVO X200 Ultra to collect images for 4 separate focal lengths (14 mm, 35 mm, 84 mm, and 200 mm). We first use PP-OCRV5 [cui2025paddleocr30technicalreport] for the rough annotation, then we manually filter the images and annotations. During the filtering and annotation, each image undergoes the following rules:
@@ 10 Dataset Collection
11 More Implementation Details
@@
Dataset Settings. For training, we combine a synthetic dataset with real paired datasets (Real-CE [ma2023benchmark] and UZ-ST). Our synthetic dataset builds upon LSDIR [li2023lsdir], containing 27,000 triplets $(x_{H},x_{L},x_{m})$. We render text on the GT of LSDIR and the corresponding text mask using the LBTS [tang2023scene], then apply Real-ESRGAN degradation wang2021real to generate LR images. Since there are misaligned image pairs in Real-CE [hu2025text, zhang2024diffusion] and text lines not annotated, we filter out misaligned image pairs and reannotate the images manually. In the end, we obtain 337 training pairs and 188 testing pairs from the Real-CE dataset. Images under 13mm and 52mm focal lengths are considered as LR $x_{L}$ and GT $x_{H}$, respectively. Following hu2025text, we use SAM-TS [ye2024hi] to obtain the text mask $x_{m}$ from $x_{H}$. UZ-ST triplets $(x_{H},x_{L},x_{m})$ are processed identically. For stage 1 training, we use cropped text regions from the dataset, while for stage 2, we use full images. Following [DBLP:conf/iclr/TuoXHGX24], we use PP-OCRv3 [DBLP:journals/corr/abs-2206-03001] to extract strings from the text region of predicted SR for evaluation of OCR-A.
@@
Training Details. For stage 1, we build our model based on the IDM baseline of DiffTSR [zhang2024diffusion]. We set $\lambda_{td}$, $\lambda_{Seg}$, $\lambda_{Focal}$, and $\lambda_{Dice}$ as $1$, $0.1$, 2020, and $1$ respectively. In stage 2, our model is built upon Stable Diffusion 3.5 medium [esser2024scaling] and follows a similar tile-based inference strategy to TADiSR [hu2025text] and SUPIR [yu2024scaling]. The timestep $t$ is set to $150$. The conditioning scale of the ControlNet is set to 1.0. $\lambda_{l2}$, $\lambda_{LPIPS}$, and $\lambda_{edge}$ are set to $1$, 55, 100100 when using the synthetic dataset for pretraining. Then, $\lambda_{edge}$ is set to 0 and the remaining coefficients are kept unchanged when training on real paired datasets. We train both stages of the model using the AdamW [loshchilov2017decoupled] optimizer and set the learning rate to $5\times 10^{-5}$ and $5\times 10^{-6}$ for stages 1 and 2 separately. All experiments are conducted on NVIDIA H20 GPUs. For stage 1, we train the model on synthetic and real datasets for 8 epochs, then finetune it on synthetic data for 2 epochs. For stage 2, we first pretrain the model on the synthetic dataset for 50 epochs. Then we train on the real paired datasets for 50 epochs.
12 More details of UZ-ST
@@
In Tab. 7, we provide detailed statistics on the composition of the UZ-ST dataset. Additionally, in Fig. 10, we present some example images from the dataset.
Figure 6: Alignment method comparison.
13 Alignment Comparison
We compare our proposed Cascade Coarse-to-Fine alignment strategy with the alignment method proposed by RealSR [cai2019toward] and single-time alignment using SIFT. During the comparison, we implement our strategy using SIFT and utilize the raw images from the 35mm dataset of our proposed UZ-ST dataset for evaluation. For metrics, we adopt PSNR, Mean Squared Error (MSE), SSIM [wang2019textsr], Normalized Cross-Correlation (NCC), and Average Keypoint Distance (AKD) to assess the alignment quality of the aligned image.
As shown in Tab. 8 and Fig. 6, pixel-based optimization methods break when handling heavily degraded images due to significant loss of details. Keypoint-based alignment like SIFT suffers from the lack of valid keypoint for matching. While our alignment method outperforms all other methods and achieves the best results, this is due to the progressive cascade alignment strategy, which breaks the hard alignment process into simpler tasks, facilitating the bridging between images taken under different focal lengths.
14 Additional results
We provide detailed quantitative results of UZ-ST in Tab. 11. Additional qualitative results are provided in Fig. 7 and Fig. 8. Fig. 9 also shows the ability of our method in handling other languages like Japanese and Korean.
We also conduct additional ablation experiment on how the randomness of OCR output affect our method. As shown in Tab. 10, the text accuracy gains as the randomness of the OCR output declines. However, even under complete random OCR prediction, our method still manage to outperform TADiSR thanks to the LR condition still guiding the stage 1 to generate text structures. This proves that the reliance of our method on the OCR model is limited.
15 Efficiency Analysis
@@
As shown in Tab. 9, stage 1 is a standard diffusion process that takes multiple steps in inference, the efficiency of our model may be suboptimal compared to one-step methods. However, our method achieves state-of-the-art performance in OCR-A, which cannot be easily obtained by extending inference time.
@@
Figure 8: Additional qualitative results of TIGER.
Figure 9: Demonstration of our method under different languages. (Zoom in for more details.)
Figure 10: Detailed examples of UZ-ST.